OpenStudy (vortish):

factor 10x^3(x-4)-5x(x-4)

OpenStudy (abb0t):

\[a^b(u^t) = au^{b+t}\]

OpenStudy (vortish):

that is no help thanks for trying

OpenStudy (anonymous):

@abb0t Doesn't it have to be the case that \(a=u\)?

OpenStudy (abb0t):

ok :(

OpenStudy (vortish):

i do not mean to sound ungrateful abbot but using those formulas confuses me if you want to teach me then using a similar problem works better

OpenStudy (anonymous):

Try dividing both terms by \(5x\) to factor it out.

OpenStudy (vortish):

2x^3(x-4)-(x-4) sound right

jimthompson5910 (jim_thompson5910):

Similar Problem: 7x(x-5) - 2(x-5) factors to (7x - 2)(x - 5) because you can factor out the GCF x-5

OpenStudy (vortish):

jim thanks for the similar problem its starting to make more sense 10x^3(x-4)-5x(x-4) (10x^3-5x)(x-4)

jimthompson5910 (jim_thompson5910):

yep now factor 10x^3-5x

jimthompson5910 (jim_thompson5910):

you can factor that piece down further

OpenStudy (vortish):

5x^2

jimthompson5910 (jim_thompson5910):

gcf of 10 and 5 is ____

OpenStudy (vortish):

five goes in to ten and its self so the gcf is five

jimthompson5910 (jim_thompson5910):

the gcf of x^3 and x is ___

OpenStudy (vortish):

x^3

jimthompson5910 (jim_thompson5910):

replace x with a number if you have to say x = 9 x^3 = 9^3 x = 9 gcf of x^3 and x is the same as saying gcf of 9^3 and 9

OpenStudy (anonymous):

\[10x^3(x-4)-5x(x-4)\] \[5x(x-4)[2x^2-1]\]

OpenStudy (anonymous):

Just take out the 5x(x-4).

OpenStudy (vortish):

thanks azteck but jim was showing me how to get to that all you did is confuse me

jimthompson5910 (jim_thompson5910):

were you able to figure out the GCF of x^3 and x ?

OpenStudy (vortish):

no

jimthompson5910 (jim_thompson5910):

how about the gcf of 9^3 and 9?

OpenStudy (vortish):

9 the gcf is three so 9^3 would be 3^3 right

OpenStudy (anonymous):

@vortish You shouldn't be doing factorising exercises if you haven't completely understood the concept of HCF (Highest Common Factor)

jimthompson5910 (jim_thompson5910):

9^3 = 9*9*9 so the GCF of 9^3 and 9 is the same as saying the GCF of 9*9*9 and 9

jimthompson5910 (jim_thompson5910):

9*9*9 and 9 have one single '9' in common and this is the largest common factor so the GCF of 9^3 and 9 is 9

OpenStudy (vortish):

azteck pls leave all you are doing is ticking me off

OpenStudy (vortish):

that make sense Jim sense 9 does go into it self by one

jimthompson5910 (jim_thompson5910):

the GCF of 7^3 and 7 is 7 the GCF of 8^3 and 8 is 8 etc etc this works for any number, so in general, you can say the GCF of x^3 and x is x

OpenStudy (vortish):

right

jimthompson5910 (jim_thompson5910):

put this together: the GCF of 10 and 5 is 5 the GCF of x^3 and x is x turns into the GCF of 10x^3 and -5x is 5x so you can factor out the GCF to go from 10x^3 - 5x to 5x(2x^2 - 1)

jimthompson5910 (jim_thompson5910):

notice how if you distribute that 5x back in, you'll get 10x^3 - 5x again

OpenStudy (vortish):

Im sorry Jim and Abbot for my out burst but Azteck was not teaching me the correct way to get to the solution and im a 36 year old man trying to figure out algebra after twenty years of being out of school im trying to make a better life for my kids and he was just making it harder

OpenStudy (vortish):

ok Jim that makes sense

jimthompson5910 (jim_thompson5910):

so you can see that 10x^3(x-4)-5x(x-4) becomes (10x^3-5x)(x-4) which further completely factors to 5x(2x^2-1)(x-4)

OpenStudy (vortish):

right thanks for the help Jim and Abbot thank you for you effort in explaining this problem to me